Grade 11Mechanics

A projectile has the same range for two angles of projection. If the time of flight in two cases are t1 and t2 then(a) the range of the projectile is
  1. 1/2gt1t2
  2. 1/4 gt1t2
  3. gt1t2
  4. 1/8gt1t2
b)the velocity of projection of projectile is
  1. 1/2g(t1+t2)
  2. 1/4g(t1+t2)
  3. 1/2g(t21+t22)1/2
  4. 1/2g(t12+t22)

Poulami
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
A Projectile has Same range for 2 diff. angles  R1=R2
THIS IMPLIES THAT
Angle of projectile 1 +Angle of Projectile 2 =90(Simple rule Can be proved if required)
Let  a + b = 90
a=90-b
sin a =cos b
T1 = 2 u sin a/g ;;  T1 =2 u cos b/g ; ;cos b=T1 .g /2u
T2= 2 u sin b/g ;;  sin b=T2.g/2u
R2 = u*u*sin2b/g
=u*u*2sin b cos b/g
SUBSTITUTE THE VALUES
=u*u*2*T2.g.T1.g/g.2u.2u
=T2*T1*g/2
(WE HAVE ASSUMED THAT THE PROJECTILES HAVE SAME INITIAL VELOCITY
AND ARE PROJECTED FROM SAME POINTS SIMULTANEOUSLY)
Regards
Arun (askIITians forum expert)